13. An insurance company sets a premium of a $5 term Insurance Policy of $1000 a 45-year old man will die during the next year is 0.0036. The probablitn distribution of the random variable X that denotes the gain to the insurance company is: for selling a l-year to a man of age 45. The probability that Random variable X $5 Probabili $5 $1000 0.0036 0.9964 6s a) What is the Expected Value (mean) or? b) If the company sells 1000 policies, what is its total profit?
Answer #1

Ans:

a)

Expected value of X=0.9965*5-0.0035*9995

E(X)=-30

b)

if 1000 policies are sold,then total profit or loss=-1000*30=-30,000